# Large Language Models Are Human-Like Internally

### v1

@@ Appendix A Cross-lingual experiments
We target FPGD data from MECO in 13 languages Siegelman et al. (2022) using five multilingual XGLMs Lin et al. (2022) (564M, 1.7B, 2.9B, 4.5B, 7.5B). We analyze the $\Delta$LL of surprisal from their internal layers. Due to the unavailability of tuned-lenses for XGLMs, only logit-lenses were used. Table 7 shows $\Delta$LLs per-layer-range, and Fugure 7 shows the relationship between paramter numbers and $\Delta$LL.
Logit-lens ($\Delta$LL) Tuned-lens ($\Delta$LL)
SPR 24.51 23.35 18.84 7.80 1.81 15.78 8.92 4.87 2.53 1.27
FPGD 22.62 26.24 25.12 15.55 5.02 16.28 14.48 11.87 9.47 5.57
SPGD 241.76 191.77 124.82 30.13 3.97 68.43 20.05 5.28 1.71 3.87
TOTAL 211.73 175.06 119.29 32.39 3.06 67.39 23.51 7.82 1.61 1.56
ELAN (125–175ms) 0.79 0.38 0.18 0.30 0.49 0.22 0.38 0.77 0.67 0.50
LAN (300–400ms) 70.32 48.69 25.78 5.05 6.14 19.94 2.74 1.72 5.02 8.76
N400 (300–500ms) 57.45 33.30 14.01 12.89 32.26 11.31 6.12 16.19 29.49 37.11
EPNP (400–600ms) 78.22 58.65 35.25 7.76 3.09 29.31 6.67 1.29 1.56 4.45
P600 (500–700ms) 69.52 50.43 29.25 6.37 7.39 17.70 3.30 1.80 5.83 11.36
PNP (600–700ms) 60.47 46.76 29.58 7.32 1.56 25.38 8.09 2.30 0.73 1.61
Features coef std err t P>|>|> |t|||| [0.025 0.975]
Intercept 0.0204 0.001 28.757 0.000 0.019 0.022
stimuli[T.Federmeier+, 2007] 0.0025 0.001 3.945 0.000 0.001 0.004
stimuli[T.Fillers] 0.0011 0.000 2.281 0.023 0.000 0.002
stimuli[T.Hubbard+, 2019] -0.0109 0.001 -17.559 $0$0.012 -0.010
stimuli[T.Michaelov+, 2024] -0.0096 0.001 -15.458 $0$0.011 -0.008
stimuli[T.NS] -0.0038 0.001 -6.927 $0$0.005 -0.003
stimuli[T.S&F,2022] -0.0105 0.001 -16.822 $0$0.012 -0.009
stimuli[T.Szewczyk+, 2022] -0.0057 0.001 -9.240 $0$0.007 -0.005
stimuli[T.UCL] 0.0054 0.000 11.807 0.000 0.004 0.006
stimuli[T.W&F,2012] -0.0109 0.001 -17.593 $0$0.012 -0.010
stimuli[T.ZuCO] -0.0007 $0$1.450 0.147 -0.002 0.000
model[T.gpt2-large] 0.0010 0.001 1.633 0.103 -0.000 0.002
model[T.gpt2-medium] -0.0004 0.001 -0.542 0.588 -0.002 0.001
model[T.gpt2-xl] 0.0013 0.001 2.199 0.028 0.000 0.002
model[T.opt-1.3b] 0.0017 0.001 2.544 0.011 0.000 0.003
model[T.opt-125m] 0.0015 0.001 1.964 0.049 3.21e-06 0.003
model[T.opt-13b] 0.0017 0.001 2.505 0.012 0.000 0.003
model[T.opt-2.7b] 0.0019 0.001 2.735 0.006 0.001 0.003
model[T.opt-30b] 0.0011 0.001 1.644 0.100 -0.000 0.002
model[T.opt-6.7b] 0.0017 0.001 2.662 0.008 0.000 0.003
model[T.opt-66b] 0.0019 0.001 2.981 0.003 0.001 0.003
model[T.pythia-1.4b-deduped] -1.555e-05 0.001 -0.024 0.981 -0.001 0.001
model[T.pythia-12b-deduped] 0.0012 0.001 1.993 0.046 2.03e-05 0.002
model[T.pythia-14m] -0.0033 0.001 -2.714 0.007 -0.006 -0.001
model[T.pythia-160m-deduped] -0.0026 0.001 -3.468 0.001 -0.004 -0.001
model[T.pythia-1b-deduped] 8.195e-06 0.001 0.010 0.992 -0.002 0.002
model[T.pythia-1b-deduped-v0] 0.0012 0.002 0.737 0.461 -0.002 0.004
model[T.pythia-2.8b-deduped] 0.0005 0.001 0.814 0.416 -0.001 0.002
model[T.pythia-31m] -0.0031 0.001 -2.567 0.010 -0.005 -0.001
model[T.pythia-410m-deduped] -0.0003 0.001 -0.486 0.627 -0.002 0.001
model[T.pythia-6.9b-deduped] 0.0007 0.001 1.173 0.241 -0.000 0.002
model[T.pythia-70m-deduped] -0.0027 0.001 -3.060 0.002 -0.004 -0.001
measure[T.MAZE] -0.0261 0.001 -39.073 $0$0.027 -0.025
measure[T.N400] -0.0124 0.001 -24.251 $0$0.013 -0.011
measure[T.SPR] -0.0030 0.001 -5.303 $0$0.004 -0.002
method[T.tuned-lens] 7.998e-05 0.000 0.389 0.697 -0.000 0.000
normalized_layer -0.0124 0.001 -20.423 $0$0.014 -0.011
measure[T.MAZE]:normalized_layer 0.0782 0.001 74.603 0.000 0.076 0.080
measure[T.N400]:normalized_layer 0.0170 0.001 22.907 0.000 0.016 0.018
measure[T.SPR]:normalized_layer 0.0009 0.001 0.978 0.328 -0.001 0.003
Table 5: Coeffecients of the regression results in Figure 4(a). The coefficients for the interaction terms listed in the last three lines are of interst.
Features coef std err t P>|>|> |t|||| [0.025 0.975]
Intercept 0.0054 0.000 20.302 0.000 0.005 0.006
stimuli[T.Federmeier+, 2007] 0.0366 0.000 137.401 0.000 0.036 0.037
stimuli[T.Fillers] 0.0007 0.000 3.936 0.000 0.000 0.001
stimuli[T.Hubbard+, 2019] 0.0232 0.000 87.206 0.000 0.023 0.024
stimuli[T.Michaelov+, 2024] 0.0245 0.000 92.110 0.000 0.024 0.025
stimuli[T.NS] -0.0014 $0$6.827 $0$0.002 -0.001
stimuli[T.S&F,2022] 0.0237 0.000 88.926 0.000 0.023 0.024
stimuli[T.Szewczyk+, 2022] 0.0284 0.000 106.623 0.000 0.028 0.029
stimuli[T.UCL] 0.0020 0.000 11.513 0.000 0.002 0.002
stimuli[T.W&F,2012] 0.0232 0.000 87.126 0.000 0.023 0.024
stimuli[T.ZuCO] 0.0232 0.000 123.112 0.000 0.023 0.024
model[T.gpt2-large] 0.0004 0.000 1.702 0.089 -5.95e-05 0.001
model[T.gpt2-medium] -0.0002 $0$0.561 0.575 -0.001 0.000
model[T.gpt2-xl] 0.0008 0.000 3.477 0.001 0.000 0.001
model[T.opt-1.3b] 0.0008 0.000 3.115 0.002 0.000 0.001
model[T.opt-125m] 0.0003 0.000 1.231 0.218 -0.000 0.001
model[T.opt-13b] 0.0008 0.000 2.952 0.003 0.000 0.001
model[T.opt-2.7b] 0.0009 0.000 3.425 0.001 0.000 0.001
model[T.opt-30b] 0.0006 0.000 2.412 0.016 0.000 0.001
model[T.opt-6.7b] 0.0005 0.000 2.184 0.029 5.24e-05 0.001
model[T.opt-66b] 0.0009 0.000 3.748 0.000 0.000 0.001
model[T.pythia-1.4b-deduped] 0.0004 0.000 1.449 0.147 -0.000 0.001
model[T.pythia-12b-deduped] 0.0010 0.000 4.172 0.000 0.001 0.001
model[T.pythia-14m] -0.0016 $0$3.566 $0$0.003 -0.001
model[T.pythia-160m-deduped] -0.0003 $0$1.082 0.279 -0.001 0.000
model[T.pythia-1b-deduped] 1.503e-05 0.000 0.047 0.962 -0.001 0.001
model[T.pythia-1b-deduped-v0] 0.0008 0.001 1.385 0.166 -0.000 0.002
model[T.pythia-2.8b-deduped] 0.0003 0.000 1.218 0.223 -0.000 0.001
model[T.pythia-31m] -0.0014 $0$2.983 0.003 -0.002 -0.000
model[T.pythia-410m-deduped] -8.252e-05 $0$0.337 0.736 -0.001 0.000
model[T.pythia-6.9b-deduped] 0.0008 0.000 3.247 0.001 0.000 0.001
model[T.pythia-70m-deduped] -0.0008 $0$2.471 0.013 -0.001 -0.000
measure[T.MAZE] -0.0046 $0$18.669 $0$0.005 -0.004
measure[T.N400] -0.0318 $0$135.362 $0$0.032 -0.031
measure[T.SPR] -0.0053 $0$24.991 $0$0.006 -0.005
method[T.tuned-lens] -0.0009 7.74e-05 -11.156 $0$0.001 -0.001
normalized_layer -0.0079 $0$35.798 $0$0.008 -0.008
measure[T.MAZE]:normalized_layer 0.0124 0.000 32.324 0.000 0.012 0.013
measure[T.N400]:normalized_layer 0.0145 0.000 52.124 0.000 0.014 0.015
measure[T.SPR]:normalized_layer 0.0094 0.000 27.774 0.000 0.009 0.010
Table 6: Coeffecients of the regression results in Figure 4(b) (only sentence/clause-final tokens are targeted).
Logit-lens ($\Delta$LL)
Lang. 0-0.2 0.2-0.4 0.4-0.6 0.6-0.8 0.8-1
Du 5.62 5.58 7.66 11.60 20.77
Ee 2.16 2.34 2.74 3.37 3.26
En 4.51 12.49 21.43 23.53 20.04
Fi 7.61 8.72 10.44 15.59 16.25
Ge 18.20 22.83 28.66 34.41 30.57
Gr 15.05 19.75 21.20 21.61 18.19
He 2.12 1.95 1.25 1.96 1.98
It 16.55 18.85 24.19 29.50 32.42
Ko 3.09 4.02 4.74 5.73 5.96
No 22.80 21.56 23.43 25.29 27.27
Ru 0.51 0.47 0.37 0.28 0.33
Sp 13.58 16.29 18.68 20.00 16.83
Tr 2.99 3.91 6.03 10.97 15.44
Table 7: Results in MECO. $\Delta$LLs are multiplied by 1000 for brevity.
Figure 7: Scaling effect between $\Delta$LL and parameter counts in MECO. The grey lines are results replying on the last layer’s $\Delta$LLs and, the red lines rely on the best internal layers’ $\Delta$LLs. In four languages (Fi, Gr, Ru, Sp) out of the targeted 13 languages, the scaling effect flipped from negative to positive, and we never observed the case of flipping from positive to negative.
Features coef std err t P>|>|> |t|||| [0.025 0.975]
@@
Table 8: The results of regression model to predict the per-token squared residual errors. Note that word frequency and length are confounded (Pearson’s $r=-0.7$), and once the length factor is excluded, the coefficient for the word frequency feature becomes significantly negative and has a high t-score. That is, a large decrease in regression error is generally associated with infrequent, long words.

### v2

@@ Appendix C Details on cross-lingual experiments
We target FPGD data from MECO in 13 languages Siegelman et al. (2022) using five multilingual XGLMs Lin et al. (2022) (564M, 1.7B, 2.9B, 4.5B, 7.5B). For the German part, we used monolingual German models in Table 4. We analyze the $\Delta$LL of surprisal from their internal layers. Due to the unavailability of tuned-lenses for XGLMs, only logit-lenses were used. Figure 5 shows the relationship between paramter numbers and $\Delta$LL in each language.
Model URL #params
GPT2 https://huggingface.co/gpt2 117M
GPT2-medium https://huggingface.co/gpt2-medium 345M
GPT2-large https://huggingface.co/gpt2-large 774M
GPT2-xl https://huggingface.co/gpt2-xl 1B
OPT-125m https://huggingface.co/facebook/opt-125m 125M
OPT-1.3b https://huggingface.co/facebook/opt-1.3b 1.3B
OPT-2.7b https://huggingface.co/facebook/opt-2.7b 2.7B
OPT-6.7b https://huggingface.co/facebook/opt-6.7b 6.7B
OPT-13b https://huggingface.co/facebook/opt-13b 13B
OPT-30b https://huggingface.co/facebook/opt-30b 30B
OPT-66b https://huggingface.co/facebook/opt-66b 66B
Pythia-14m-deduped https://huggingface.co/EleutherAI/pythia-14m-deduped 14M
Pythia-31m-deduped https://huggingface.co/EleutherAI/pythia-31m-deduped 31M
Pythia-70m-deduped https://huggingface.co/EleutherAI/pythia-70m-deduped 70M
Pythia-160m-deduped https://huggingface.co/EleutherAI/pythia-160m-deduped 160M
Pythia-410m-deduped https://huggingface.co/EleutherAI/pythia-410m-deduped 410M
Pythia-1b-deduped https://huggingface.co/EleutherAI/pythia-1b-deduped 1B
Pythia-1.4b-deduped https://huggingface.co/EleutherAI/pythia-1.4b-deduped 1.4B
Pythia-2.8b-deduped https://huggingface.co/EleutherAI/pythia-2.8b-deduped 2.8B
Pythia-6.9b-deduped https://huggingface.co/EleutherAI/pythia-6.9b-deduped 6.9B
Pythia-12b-deduped https://huggingface.co/EleutherAI/pythia-12b-deduped 12B
Llama-3.1-8B https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct 8B
Llama-3.1-70B https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct 70B
Qwen2.5-0.5B https://huggingface.co/Qwen/Qwen2.5-0.5B 500M
Qwen2.5-1.5B https://huggingface.co/Qwen/Qwen2.5-1.5B 1.5B
Qwen2.5-3B https://huggingface.co/Qwen/Qwen2.5-3B 3B
Qwen2.5-7B https://huggingface.co/Qwen/Qwen2.5-7B 7B
Qwen2.5-14B https://huggingface.co/Qwen/Qwen2.5-14B 14B
Qwen2.5-32B https://huggingface.co/Qwen/Qwen2.5-32B 32B
Qwen2.5-72B https://huggingface.co/Qwen/Qwen2.5-72B 72B
XGLM-564M https://huggingface.co/facebook/xglm-564M 564M
XGLM-1.7B https://huggingface.co/facebook/xglm-1.7B 1.7B
XGLM-2.9B https://huggingface.co/facebook/xglm-2.9B 2.9B
XGLM-4.5B https://huggingface.co/facebook/xglm-4.5B 4.5B
XGLM-7.5B https://huggingface.co/facebook/xglm-7.5B 7.5B
German LMs https://huggingface.co/benjamin/gerpt2 124M
https://huggingface.co/benjamin/gerpt2-large 774M
https://huggingface.co/malteos/gpt2-xl-wechsel-german 1.5B
Table 4: LM details
coef std err t P>|>|> |t||| [0.025 0.975]
Intercept 0.0125 0.001 20.725 0.000 0.011 0.014
stimuli[T.Federmeier+, 2007] 0.0040 0.001 7.520 0.000 0.003 0.005
stimuli[T.Fillers] 0.0002 0.000 0.570 0.569 -0.001 0.001
stimuli[T.Hubbard+, 2019] -0.0079 0.001 -14.755 $0$0.009 -0.007
stimuli[T.Michaelov+, 2024] -0.0068 0.001 -12.666 $0$0.008 -0.006
stimuli[T.NS] -0.0069 $0$14.314 $0$0.008 -0.006
stimuli[T.S&F,2022] -0.0075 0.001 -13.955 $0$0.009 -0.006
stimuli[T.Szewczyk+, 2022] -0.0033 0.001 -6.076 $0$0.004 -0.002
stimuli[T.UCL] 0.0090 0.000 22.701 0.000 0.008 0.010
stimuli[T.W&F,2012] -0.0079 0.001 -14.816 $0$0.009 -0.007
stimuli[T.ZuCO] 0.0036 0.000 8.735 0.000 0.003 0.004
model[T.Llama-3.1-8B] 0.0010 0.001 1.707 0.088 -0.000 0.002
model[T.Qwen2.5-0.5B] -0.0004 0.001 -0.689 0.491 -0.002 0.001
model[T.Qwen2.5-1.5B] -0.0007 0.001 -1.101 0.271 -0.002 0.001
model[T.Qwen2.5-14B] -0.0004 0.001 -0.822 0.411 -0.001 0.001
model[T.Qwen2.5-32B] 0.0002 0.000 0.473 0.637 -0.001 0.001
model[T.Qwen2.5-3B] -0.0012 0.001 -2.129 0.033 -0.002 -9.41e-05
model[T.Qwen2.5-72B] 0.0003 0.000 0.731 0.465 -0.001 0.001
model[T.Qwen2.5-7B] -0.0003 0.001 -0.481 0.630 -0.001 0.001
model[T.gpt2] 0.0010 0.001 1.568 0.117 -0.000 0.002
model[T.gpt2-large] 0.0020 0.000 4.266 0.000 0.001 0.003
model[T.gpt2-medium] 0.0008 0.001 1.242 0.214 -0.000 0.002
model[T.gpt2-xl] 0.0023 0.000 5.367 0.000 0.001 0.003
model[T.opt-1.3b] 0.0026 0.001 5.084 0.000 0.002 0.004
model[T.opt-125m] 0.0024 0.001 3.778 0.000 0.001 0.004
model[T.opt-13b] 0.0028 0.001 5.301 0.000 0.002 0.004
model[T.opt-2.7b] 0.0028 0.001 4.902 0.000 0.002 0.004
model[T.opt-30b] 0.0023 0.001 4.513 0.000 0.001 0.003
model[T.opt-6.7b] 0.0027 0.000 5.639 0.000 0.002 0.004
model[T.opt-66b] 0.0027 0.000 5.721 0.000 0.002 0.004
model[T.pythia-1.4b-deduped] 0.0010 0.001 1.992 0.046 1.65e-05 0.002
model[T.pythia-12b-deduped] 0.0023 0.000 4.920 0.000 0.001 0.003
model[T.pythia-14m] -0.0017 0.001 -1.508 0.132 -0.004 0.000
model[T.pythia-160m-deduped] -0.0015 0.001 -2.352 0.019 -0.003 -0.000
model[T.pythia-1b-deduped] 0.0010 0.001 1.367 0.172 -0.000 0.002
model[T.pythia-1b-deduped-v0] 0.0040 0.002 2.415 0.016 0.001 0.007
model[T.pythia-2.8b-deduped] 0.0016 0.000 3.349 0.001 0.001 0.003
model[T.pythia-31m] -0.0015 0.001 -1.391 0.164 -0.004 0.001
model[T.pythia-410m-deduped] 0.0008 0.001 1.476 0.140 -0.000 0.002
model[T.pythia-6.9b-deduped] 0.0018 0.000 3.826 0.000 0.001 0.003
model[T.pythia-70m-deduped] -0.0016 0.001 -1.911 0.056 -0.003 4.05e-05
measure[T.FPGD] 0.0053 0.000 10.675 0.000 0.004 0.006
measure[T.N400] -0.0069 0.001 -13.720 $0$0.008 -0.006
measure[T.MAZE] -0.0160 0.001 -28.315 $0$0.017 -0.015
method[T.tuned-lens] 0.0003 0.000 1.365 0.172 -0.000 0.001
normalized_layer -0.0076 0.001 -12.601 $0$0.009 -0.006
measure[T.FPGD]:layer -0.0034 0.001 -4.242 $0$0.005 -0.002
measure[T.N400]:layer 0.0100 0.001 14.084 0.000 0.009 0.011
measure[T.MAZE]:layer 0.0593 0.001 61.971 0.000 0.057 0.061
(a) All data
coef std err t P>|>|> |t||| [0.025 0.975]
Intercept -0.0002 $0$0.971 0.331 -0.001 0.000
stimuli[T.Federmeier+, 2007] 0.0394 0.000 168.408 0.000 0.039 0.040
stimuli[T.Fillers] 0.0011 0.000 7.379 0.000 0.001 0.001
stimuli[T.Hubbard+, 2019] 0.0275 0.000 117.432 0.000 0.027 0.028
stimuli[T.Michaelov+, 2024] 0.0286 0.000 122.213 0.000 0.028 0.029
stimuli[T.NS] -0.0011 $0$6.066 $0$0.001 -0.001
stimuli[T.S&F,2022] 0.0279 0.000 119.263 0.000 0.027 0.028
stimuli[T.Szewczyk+, 2022] 0.0321 0.000 137.295 0.000 0.032 0.033
stimuli[T.UCL] 0.0029 0.000 19.160 0.000 0.003 0.003
stimuli[T.W&F,2012] 0.0275 0.000 117.292 0.000 0.027 0.028
stimuli[T.ZuCO] 0.0274 0.000 165.961 0.000 0.027 0.028
model[T.Llama-3.1-8B] 6.679e-05 0.000 0.299 0.765 -0.000 0.001
model[T.Qwen2.5-0.5B] -0.0002 $0$0.667 0.505 -0.001 0.000
model[T.Qwen2.5-1.5B] 9.912e-05 0.000 0.424 0.672 -0.000 0.001
model[T.Qwen2.5-14B] 3.475e-05 0.000 0.178 0.859 -0.000 0.000
model[T.Qwen2.5-32B] -2.221e-05 $0$0.123 0.902 -0.000 0.000
model[T.Qwen2.5-3B] -0.0003 $0$1.455 0.146 -0.001 0.000
model[T.Qwen2.5-72B] -0.0004 $0$2.508 0.012 -0.001 -9.3e-05
model[T.Qwen2.5-7B] -0.0005 $0$2.061 0.039 -0.001 -2.36e-05
model[T.gpt2] -0.0003 $0$1.304 0.192 -0.001 0.000
model[T.gpt2-large] 3.558e-05 0.000 0.199 0.842 -0.000 0.000
model[T.gpt2-medium] -0.0005 $0$2.208 0.027 -0.001 -6.13e-05
model[T.gpt2-xl] 0.0005 0.000 2.718 0.007 0.000 0.001
model[T.opt-1.3b] 0.0004 0.000 1.794 0.073 -3.3e-05 0.001
model[T.opt-125m] -0.0001 $0$0.411 0.681 -0.001 0.000
model[T.opt-13b] 0.0004 0.000 1.747 0.081 -4.41e-05 0.001
model[T.opt-2.7b] 0.0005 0.000 2.341 0.019 8.51e-05 0.001
model[T.opt-30b] 0.0002 0.000 1.040 0.298 -0.000 0.001
model[T.opt-6.7b] 0.0001 0.000 0.669 0.503 -0.000 0.000
model[T.opt-66b] 0.0005 0.000 2.709 0.007 0.000 0.001
model[T.pythia-1.4b-deduped] -5.816e-05 $0$0.293 0.770 -0.000 0.000
model[T.pythia-12b-deduped] 0.0005 0.000 2.831 0.005 0.000 0.001
model[T.pythia-14m] -0.0018 $0$4.125 $0$0.003 -0.001
model[T.pythia-160m-deduped] -0.0008 $0$3.296 0.001 -0.001 -0.000
model[T.pythia-1b-deduped] -0.0004 $0$1.331 0.183 -0.001 0.000
model[T.pythia-1b-deduped-v0] 0.0002 0.001 0.294 0.769 -0.001 0.001
model[T.pythia-2.8b-deduped] -0.0001 $0$0.646 0.519 -0.000 0.000
model[T.pythia-31m] -0.0018 $0$4.200 0.000 -0.003 -0.001
model[T.pythia-410m-deduped] -0.0005 $0$2.451 0.014 -0.001 -9.76e-05
model[T.pythia-6.9b-deduped] 0.0003 0.000 1.807 0.071 -2.81e-05 0.001
model[T.pythia-70m-deduped] -0.0013 $0$4.092 $0$0.002 -0.001
measure[T.FPGD] 0.0053 0.000 27.964 0.000 0.005 0.006
measure[T.N400] -0.0301 $0$124.125 $0$0.031 -0.030
measure[T.MAZE] 0.0007 0.000 3.082 0.002 0.000 0.001
method[T.tuned-lens] -0.0003 8.25e-05 -3.448 0.001 -$0$0.000
normalized_layer 0.0015 0.000 6.663 0.000 0.001 0.002
measure[T.FPGD]:layer -0.0088 $0$29.286 $0$0.009 -0.008
measure[T.N400]:layer 0.0046 0.000 16.990 0.000 0.004 0.005
measure[T.MAZE]:layer 0.0030 0.000 8.430 0.000 0.002 0.004
(b) Only with sentence/clause-final tokens
Table 5: Obtained coefficients in Section˜6.2
Features coef std err t P>|>|> |t||| [0.025 0.975]
@@
Table 6: Obtained coefficients in Section˜6.4. Note that word frequency and length are confounded (Pearson’s $r=-0.7$), and once the length factor is excluded, the coefficient for the word frequency feature becomes significantly negative and has a high t-score. That is, a large decrease in regression error is generally associated with infrequent, long words.
Logit-lens ($\Delta$LL) Tuned-lens ($\Delta$LL)
SPR 24.51 23.35 18.84 7.80 1.81 15.78 8.92 4.87 2.53 1.27
FPGD 22.11 23.39 22.83 15.77 6.53 16.28 14.48 11.87 9.47 5.57
SPGD 228.29 189.49 144.70 72.48 21.47 68.43 20.05 5.28 1.71 3.87
TOTAL 201.36 171.44 134.24 68.64 19.82 67.39 23.51 7.82 1.61 1.56
ELAN (125–175ms) 0.75 0.48 0.31 0.36 0.34 0.22 0.38 0.77 0.67 0.50
LAN (300–400ms) 67.55 51.86 35.96 17.45 8.17 19.94 2.74 1.72 5.02 8.76
N400 (300–500ms) 56.86 38.04 22.77 16.07 22.58 11.31 6.12 16.19 29.49 37.11
EPNP (400–600ms) 75.94 61.74 46.02 22.85 7.84 29.31 6.67 1.29 1.56 4.45
P600 (500–700ms) 69.02 55.91 41.88 22.60 10.37 17.70 3.30 1.80 5.83 11.36
PNP (600–700ms) 59.75 50.83 39.67 21.40 7.26 25.38 8.09 2.30 0.73 1.61
(a) Logit-lens
(b) Tuned-lens
[…]